OpenStudy (anonymous):

Let f(x)=x^2-1 and g(x)=2x-3 Find (g * f)(3)

OpenStudy (anonymous):

(g * f)(x)=g(x)*f(x)=(2x-3)(x^2-1) just put x=3 to get your answer

OpenStudy (anonymous):

Okay thanks. That was way easier than i thought.